A ball is dropped from a height of 3 m. After each bounce, it rises to 85% of its previous height. What is the total distance that the ball has travelled when it comes to rest?


1
Expert's answer
2021-09-27T15:39:35-0400

We have the infinite geometric series: "a=2(3\cdot0.85), r=0.85."


"S=\dfrac{a}{1-r}=\dfrac{6\cdot0.85}{1-0.85}=34"

The total distance that the ball has travelled when it comes to rest is


"3+34=37(m)"